OpenStudy (anonymous):

Please Help! A store sold 39 sweatshirts. White ones cost $11.95 and yellow cost $13.50.I $497.05 worth of shirts were sold. How many of each were sold?

OpenStudy (anonymous):

11.95x +13.5y = 497.05 would be your equation for the prices x + y = 39 to represent the total number of sweatshirts sold x represents number of white, and y represents number of yellow. Then you solve the equations to find the answer.

OpenStudy (anonymous):

Ok i did 11.95+13.50=497.05 subtracted 13.50 from 497.05 I got 483.55 from there do I divide that by 11.95?